how to convert quintal number 13242 to octal


1
Expert's answer
2021-02-01T19:02:18-0500

In the quintal (base 5) system we have:


132425=1×54+3×53+2×52+4×51+2×5013242_5=1\times5^4+3\times5^3+2\times5^2+4\times5^1+2\times5^0

=1072(in decimal)=1072(in\ decimal)

1072=2×83+0×82+6×81+0×801072=2\times8^3+0\times8^2+6\times8^1+0\times8^0

Therefore


13242(quintal)=2060(octal)13242(quintal)=2060(octal)